# Transaction Fee Market Design for Parallel Execution

### v1

@@ 6. Towards a Fee Market for Parallel Execution
Armed with an understanding of the achievable trade-offs between desirable properties in a GCM — specifically, the impossibility of satisfying all properties in a single mechanism — and our analysis of various candidate mechanisms and their properties, we propose the weighted area GCM for implementation in a fee market for parallel execution. Given that the weighted area GCM satisfies Easy Gas Estimation, it achieves the most additional properties one could hope for in a non-constant GCM (see Theorem 3.4).
@@
Satisfying Easy Gas Estimation ensures that each transaction consumes a fixed amount of gas, regardless of other transactions in the same block. From the perspective of currently deployed TFMs (e.g., EIP-1559), which process transactions with fixed sizes, nothing changes. Thus, the existing properties of these mechanisms remain intact. At a high level, the key properties we strive for in a TFM are incentive compatibility for both block producers and users, welfare optimality, and collusion resistance. However, no TFM can achieve all these properties simultaneously (Chung and Shi, 2023; Chung et al., 2024; Gafni and Yaish, 2024a). Importantly, when composing the weighted area GCM with a TFM of choice, the level of sophistication required from users in their bidding strategy does not increase—unlike in currently deployed fee markets for parallel execution (Lostin, 2025; Mueller, 2025).
We next further outline some additional steps that need to be taken to implement the weighted area GCM in practice:
Setting Weights.
The weighted area GCM uses weights to reflect resource demand, i.e., the higher the demand for a resource, the higher its associated weight. As a result, transactions that are harder to parallelize due to execution on high-demand resources will incur higher gas costs. An open question remains on how to determine the relative weights of these resources. One potential approach is to use an adjustment mechanism similar to those employed in EIP-1559 (Buterin et al., 2024), the blob fee market (Buterin et al., 2022) (an additional Ethereum fee market for data availability), or other proposals aimed at addressing limitations in current demand adjustment mechanisms (Azouvi et al., 2023).
Integration with Other Gas Components.
Importantly, the weighted area GCM we propose focuses solely on the execution component, whereas gas, in practice, accounts for multiple factors, such as data bandwidth and storage. Moving toward real-world deployment could take two approaches:
— One option is integrating the new execution gas measure into Ethereum’s existing framework, which already assigns gas weights to different transaction components (e.g., storage and data bandwidth). These weights are then combined into a single value representing transaction size. To incorporate the weighted area GCM, one would need to determine the relative contribution of execution compared to other gas components and adjust the scaling factors accordingly. With the new weighted area GCM, this task becomes even more challenging, as it is difficult to establish a single value that remains fixed over several years. Specifically, determining the relative cost of reduced parallelization potential from a transaction, compared to the cost of a byte of bandwidth, presents a significant challenge.
— A more principled approach would be to integrate the new execution gas measure into a multi-dimensional fee market, where execution is one component alongside storage, data bandwidth, and other factors. The Ethereum community is actively discussing the transition to a multi-dimensional fee market (Buterin, 2022), making this a timely opportunity to incorporate the new execution gas measure. The execution gas component we propose can itself be understood as a multi-dimensional fee market, where each resource represents a distinct dimension with its own posted price, dynamically adjusted based on demand, similarly to the previously outlined approach for weights. These resource dimensions could then be integrated as top-level dimensions within the full multi-dimensional fee market. For example, one dimension might represent bandwidth, while another could correspond to execution on resource $r_{5}$. Beyond the general advantages of a multi-dimensional fee market, this approach is perhaps also the most principled and effective method to integrate the weighted area GCM into a fee market.
7. Related Work
7.1. Transaction Fee Mechanisms
@@
There is extensive research on blockchain fee markets, with a particular focus on Ethereum and Bitcoin. Early studies primarily examined Bitcoin, exploring monopolistic pricing mechanisms (Lavi et al., 2022; Yao, 2020). More recent contributions to this field include (Nisan, 2023; Gafni and Yaish, 2022; Penna and Schneider, 2024). Unlike these works, our study concerns measuring resource usage on a blockchain with client-side parallel execution, rather than focusing on pricing.
@@
The TFM design framework was introduced by Roughgarden (Roughgarden, 2020, 2024). Roughgarden’s analysis of the EIP-1559 mechanism (Buterin et al., 2024) initiated an active line of research on TFMs. Chung and Shi (Chung and Shi, 2023) demonstrated that no TFM can be ideal — meaning it cannot simultaneously be incentive-compatible for users and block producers while also being resistant to collusion between the two. This conclusion holds even for weaker definitions of collusion resilience, as shown by Chung et al. (Chung et al., 2024) and Gafni and Yaish (Gafni and Yaish, 2024a). Finally, attempts to address these limitations using cryptographic techniques (Shi et al., 2022; Wu et al., 2023) have made progress in overcoming certain impossibilities, while other attempts relax the desiderata (Gafni and Yaish, 2024b). However, designing an ideal TFM still remains out of reach. While these studies examine the limitations of TFMs, our focus is on GCMs for parallel execution and how to integrate them with a TFM.
@@
A related body of work examines the dynamics of TFMs over multiple blocks, particularly focusing on the base fee in EIP-1559. Leonardos et al. (Leonardos et al., 2021, 2023) demonstrate that the stability of the base fee depends on the adjustment parameter, with short-term volatility but long-term block size stability. Reijsbergen et al. (Reijsbergen et al., 2021) suggest using an adaptive adjustment parameter to mitigate block size fluctuations, while Ferreira et al. (Ferreira et al., 2021) highlight user experience issues caused by bounded base fee oscillations. Additionally, Hougaard and Pourpouneh (Hougaard and Pourpouneh, 2023) and Azouvi et al. (Azouvi et al., 2023) reveal that the base fee can be manipulated by non-myopic miners.
@@
Given the discussion surrounding multi-dimensional fees in Ethereum (Buterin, 2022, 2024) and the deployment of EIP-4844 (Buterin et al., 2022) (a first step towards a multi-dimensional fee market on Ethereum), a recent line of work explores multi-dimensional fee markets, focusing on efficient pricing mechanisms and their optimality. This work is further refined by Diamandis et al. (Diamandis et al., 2023), who design and analyze multi-dimensional blockchain fee markets to align incentives and improve network performance. Building on this, Angeris et al. (Angeris et al., 2024) prove that such fee markets are nearly optimal, with efficiency improving over time even under adversarial conditions. Multidimensional fee markets are closely related to fee markets designed for parallel execution. In particular, in the weighted area GCM, the weights can be interpreted as fees within a multidimensional fee market. Unlike previous literature on multidimensional fee markets, we focus on parallelization, introduce desirable properties, and evaluate how various mechanisms perform.
@@
Further extensions of TFMs have emerged. Bahrani et al. (Bahrani et al., 2024) consider TFMs in the presence of maximal extractable value (MEV), i.e., value extractable by the block producer. Further, Wang et al. (Wang et al., 2024) design a fee mechanism for proof networks, whereas Bahrani et al. (Bahrani and Durvasula, 2024) introduce a transaction fee mechanism for heterogeneous computation. Our work most closely relates to the latter, but, in contrast, our chosen approach is closer to multidimensional fee markets, trading complexity for the block producer for stronger incentive compatibility for the user.
@@
Local fee markets have recently been a topic of discussion in the blockchain space (Eclipse Labs, 2024; Diamandis et al., 2024; Lostin, 2025; keyneom, 2024). The core idea is that transactions interacting with highly contested states incur higher fees, while those involving non-contested states pay lower fees. However, discussions on local fee markets have largely remained high-level, without a precise characterization of the desired properties beyond this general goal. Moreover, currently implemented local fee markets (Lostin, 2025) require significant user sophistication to set fees appropriately. In this work, we formalize the desiderata for fee markets in the context of parallel execution and identify the weighted area GCM as a promising candidate. One key advantage is its compatibility with a TFM, enabling simple fee estimation for users.
7.2. Parallel Execution
@@
Blockchain concurrency has been a focal point in an active line of research. In particular, numerous efforts have aimed to enable parallel transaction processing through speculative execution (Sergey and Hobor, 2017; Zhang and Zhang, 2018; Amiri et al., 2019; Dickerson et al., 2017; Anjana et al., 2022; Gelashvili et al., 2022; Chen et al., 2021; Saraph and Herlihy, 2019). Note that speculative execution is already deployed by multiple blockchains (Aptos Labs, 2023; Sei Labs, 2023; Monad Labs, 2023). Static analysis has also been employed to identify parallelizable transactions, though it cannot completely eliminate inherent dependencies (Pîrlea et al., 2021; Murgia et al., 2021). Similarly, Neiheiser et al. (2024) demonstrate how parallel execution can assist struggling nodes in catching up. While these works are orthogonal to ours, they highlight the overhead of parallel execution when there is no advance knowledge about a transaction’s state accesses.
@@
Further, Saraph and Herlihy (2019) and Heimbach et al. (2023a) have evaluated the parallelization potential of the Ethereum workload. The latter demonstrates that a speedup of approximately fivefold is achievable, assuming state accesses are known in advance. Additionally, Solana (Solana Labs, 2023) and Sui (Mysten Labs, 2023) already perform parallel execution with advance knowledge of state accesses. However, in practice, state accesses are not known beforehand on many blockchains such as Ethereum. There, less than 2% of transactions disclose them proactively, as shown by Heimbach et al. (2023b), due to a lack of incentives. In this work, we aim to take a step toward unlocking the parallelization potential by designing a TFM that supports parallel execution. This mechanism relies on the disclosure of state accesses as done in Solana (Solana Labs, 2023) and Sui (Mysten Labs, 2023).
8. Conclusion
@@
In this work, we addressed this gap by introducing a framework with two key components: a GCM, which measures the execution-related load a transaction imposes on the network in units of gas, and a TFM, which determines the cost associated with each unit of gas. We then formalized the desired properties for the GCM in such a fee market. After outlining the desiderata, we evaluated various mechanisms against them and identified a strong candidate through this analysis — the weighted area GCM.
@@
Acknowledgements
We thank the Robust Incentives Group at the Ethereum Foundation for supporting this work through a grant. We thank Boyan Zhou for his work on a student project, supervised by the last three authors, which helped shape this work. Additionally, we are grateful to the attendees of EthereumZüri.ch 2025 for their insightful comments and feedback on this work.
Appendix A Proofs Omitted From Section 3
@@
Proof.
Properties 1 and 2 are special cases of Property 3, corresponding to the scenarios where $t_{1}=t_{2}$ and $R_{1}=R_{2}$, respectively. Therefore, the ()($\Rightarrow$)( ⇒ ) direction holds. To prove the ()($\Leftarrow$)( ⇐ ) direction, assume Properties 1 and 2 hold and consider a set of transactions $T$ and two transactions $\textit{tx}_{1}\simeq(t_{1},R_{1})$ and $\textit{tx}_{2}\simeq(t_{2},R_{2})$, both not in $T$, such that $t_{1}\leq t_{2}$ and $R_{1}\subseteq R_{2}$. Let $\textit{tx}_{3}$ be another transaction not in $T$ such that $\textit{tx}_{3}\simeq(t_{1},R_{2})$. By first applying Property 1 and then Property 2, we obtain:
@@
This establishes the conclusion. ∎
@@
Proof.
The strict variants of Properties 1 and 2 are special cases of the strict variant of Property 3, corresponding to the scenarios where $t_{1}=t_{2}$ and $R_{1}=R_{2}$, respectively. Therefore, the ()($\Rightarrow$)( ⇒ ) direction holds. To prove the ()($\Leftarrow$)( ⇐ ) direction, assume Properties 1 and 2 hold strictly and consider a set of transactions $T$ and two transactions $\textit{tx}_{1}\simeq(t_{1},R_{1})$ and $\textit{tx}_{2}\simeq(t_{2},R_{2})$, both not in $T$, such that $t_{1}\leq t_{2}$ and $R_{1}\subseteq R_{2}$, and additionally $\textit{tx}_{1}\not\simeq\textit{tx}_{2}$ (i.e., at least one of the previous holds strictly). Let $\textit{tx}_{3}$ be another transaction not in $T$ such that $\textit{tx}_{3}\simeq(t_{1},R_{2})$. By first applying Property 1 and then Property 2 (their non-strict versions), we obtain:
(5) $\textsf{gas}_{T\cup\{\textit{tx}_{1}\}}(\textit{tx}_{1})\leq\textsf{gas}_{T\cup\{\textit{tx}_{3}\}}(\textit{tx}_{3})\leq\textsf{gas}_{T\cup\{\textit{tx}_{2}\}}(\textit{tx}_{2}).$
From this, we get that $\textsf{gas}_{T\cup{\textit{tx}_{1}}}(\textit{tx}_{1})\leq\textsf{gas}_{T\cup{\textit{tx}_{2}}}(\textit{tx}_{2})$, so it remains to rule out the equality case. Assume for a contradiction that $\textsf{gas}_{T\cup\{\textit{tx}_{1}\}}(\textit{tx}_{1})=\textsf{gas}_{T\cup\{\textit{tx}_{2}\}}(\textit{tx}_{2})$, from which the two inequalities hold with equality in Eq. 5. Because Properties 1 and 2 hold strictly, this can only be the case if $\textit{tx}_{1}\simeq\textit{tx}_{3}\simeq\textit{tx}_{2}$, contradicting the assumption that $\textit{tx}_{1}\not\simeq\textit{tx}_{2}$. ∎
Appendix B Proofs Omitted From Section 4

### v2

@@ 6 Towards a Fee Market for Parallel Execution
Armed with an understanding of the trade-offs between desirable properties in a GCM, in particular the impossibility of satisfying all properties within a single mechanism, and informed by our analysis of various candidate mechanisms and their properties, we propose two mechanisms for practical implementations of a fee market supporting parallel execution. Each represents one side of the design spectrum: one drawn from the class of mechanisms with Easy Gas Estimation, and the other from the class without it.
@@
The advantage of adopting a mechanism with Easy Gas Estimation is that satisfying this property ensures that each transaction consumes a fixed amount of gas, regardless of other transactions in the same block. From the perspective of currently deployed TFMs (e.g., EIP-1559), which process transactions with fixed sizes, nothing changes. Thus, the existing properties of these mechanisms remain intact. On a high level, the key properties we strive for in a TFM are incentive compatibility for both block producers and users, welfare optimality, and collusion resistance. However, no TFM can achieve all these properties simultaneously [17, 16, 26]. Importantly, when composing a GCM that satisfies Easy Gas Estimation with a TFM of choice, the level of sophistication required from users in their bidding strategy does not increase, unlike in currently deployed fee markets for parallel execution [39, 43].
On the other hand, foregoing Easy Gas Estimation makes it possible to price transactions according to the load they impose on the network relative to the other transactions in the block, rather than evaluating each one in isolation. Mechanisms in this class can align the total gas charged in a block with the actual execution cost of that block, enabling resource pricing at the block level rather than only at the transaction level. Recent discussions in Ethereum research have explored similar designs under the umbrella of block-level fee markets [10]. If the protocol is already moving toward block-level metering for other resources, such as data availability or storage contention, extending this approach to execution costs could enable direct efficiency gains from parallel execution that are not achievable with transaction-level pricing alone.
For the class of mechanisms with Easy Gas Estimation, the natural choice is the Weighted Area GCM. Given that it satisfies Easy Gas Estimation, it achieves the most additional properties one could hope for in a (non-constant) GCM (see Theorem˜3.4).
For the class of mechanisms without Easy Gas Estimation, we propose the Time-Proportional Makespan GCM as a concrete candidate. By allocating gas in proportion to each transaction’s execution time relative to the block’s total makespan, it provides a simple way to align gas consumption with how transactions constrain parallel execution within a block.
The decision between these approaches ultimately depends on protocol-level priorities: maintaining user-side simplicity and compatibility with existing TFMs through Easy Gas Estimation, which also makes implementation simpler by reusing current transaction-level pricing infrastructure, or pursuing block-level efficiency and more accurate resource pricing at the cost of giving up the previous benefits.
7 Related Work
7.1 Transaction Fee Mechanisms
@@
There is extensive research on blockchain fee markets, with a particular focus on Ethereum and Bitcoin. Early studies primarily examined Bitcoin, exploring monopolistic pricing mechanisms [36, 63]. More recent contributions to this field include [47, 25, 49]. Unlike these works, our study concerns measuring storage key usage on a blockchain with client-side parallel execution, rather than focusing on pricing.
@@
The TFM design framework was introduced by Roughgarden [53, 54]. Roughgarden’s analysis of the EIP-1559 mechanism [11] initiated an active line of research on TFMs. Chung and Shi [17] demonstrated that no TFM can be ideal — meaning it cannot simultaneously be incentive-compatible for users and block producers while also being resistant to collusion between the two. This conclusion holds even for weaker definitions of collusion resilience, as shown by Chung et al. [16] and Gafni and Yaish [26]. Finally, attempts to address these limitations using cryptographic techniques [58, 61] have made progress in overcoming certain impossibilities, while other attempts relax the desiderata [27]. However, designing an ideal TFM still remains out of reach. While these studies examine the limitations of TFMs, our focus is on GCMs for parallel execution and how to integrate them with a TFM.
@@
A related body of work examines the dynamics of TFMs over multiple blocks, particularly focusing on the base fee in EIP-1559. Leonardos et al. [37, 38] demonstrate that the stability of the base fee depends on the adjustment parameter, with short-term volatility but long-term block size stability. Reijsbergen et al. [52] suggest using an adaptive adjustment parameter to mitigate block size fluctuations, while Ferreira et al. [24] highlight user experience issues caused by bounded base fee oscillations. Additionally, Hougaard and Pourpouneh [31] and Azouvi et al. [5] reveal that the base fee can be manipulated by non-myopic miners.
@@
Given the discussion surrounding multi-dimensional fees in Ethereum [8, 9] and the deployment of EIP-4844 [12] (a first step towards a multi-dimensional fee market on Ethereum), a recent line of work explores multi-dimensional fee markets, focusing on efficient pricing mechanisms and their optimality. This work is further refined by Diamandis et al. [20], who design and analyze multi-dimensional blockchain fee markets to align incentives and improve network performance. Building on this, Angeris et al. [2] prove that such fee markets are nearly optimal, with efficiency improving over time even under adversarial conditions. Multidimensional fee markets are closely related to fee markets designed for parallel execution. In particular, in the weighted area GCM, the weights can be interpreted as fees within a multidimensional fee market. Unlike previous literature on multidimensional fee markets, we focus on parallelization, introduce desirable properties, and evaluate how various mechanisms perform.
@@
Further extensions of TFMs have emerged. Bahrani et al. [7] consider TFMs in the presence of maximal extractable value (MEV), i.e., value extractable by the block producer. Further, Wang et al. [60] design a fee mechanism for proof networks, whereas Bahrani et al. [6] introduce a transaction fee mechanism for heterogeneous computation. Our work most closely relates to the latter, but, in contrast, our chosen approach is closer to multidimensional fee markets, trading complexity for the block producer for stronger incentive compatibility for the user.
@@
Local fee markets have recently been a topic of discussion in the blockchain space [22, 19, 39, 33]. The core idea is that transactions interacting with highly contested states incur higher fees, while those involving non-contested states pay lower fees. However, discussions on local fee markets have largely remained high-level, without a precise characterization of the desired properties beyond this general goal. Moreover, currently implemented local fee markets [39] require significant user sophistication to set fees appropriately. In this work, we formalize the desiderata for fee markets in the context of parallel execution and identify the weighted area GCM as a promising candidate. One key advantage is its compatibility with a TFM, enabling simple fee estimation for users.
7.2 Parallel Execution
@@
Blockchain concurrency has been a focal point in an active line of research. In particular, numerous efforts have aimed to enable parallel transaction processing through speculative execution [57, 64, 1, 21, 3, 28, 15, 55]. Note that speculative execution is already deployed by multiple blockchains [4, 56, 42]. Static analysis has also been employed to identify parallelizable transactions, though it cannot completely eliminate inherent dependencies [50, 44]. Similarly, Neiheiser et al. [46] demonstrate how parallel execution can assist struggling nodes in catching up. While these works are orthogonal to ours, they highlight the overhead of parallel execution when there is no advance knowledge about a transaction’s state accesses.
@@
Further, Saraph and Herlihy [55] and Heimbach et al. [29] have evaluated the parallelization potential of the Ethereum workload. The latter demonstrates that a speedup of approximately fivefold is achievable, assuming state accesses are known in advance. Additionally, Solana [59] and Sui [45] already perform parallel execution with advance knowledge of state accesses. However, in practice, state accesses are not known beforehand on many blockchains such as Ethereum. There, less than 2% of transactions disclose them proactively, as shown by Heimbach et al. [30], due to a lack of incentives. In this work, we aim to take a step toward unlocking the parallelization potential by designing a TFM that supports parallel execution. This mechanism relies on the disclosure of state accesses as done in Solana [59] and Sui [45].
7.3 Cloud Computing Pricing and Parallel Resource Allocation
While our primary focus is on transaction fee mechanisms for blockchain networks, parallels can be drawn to resource allocation and pricing in cloud computing. Several papers study pricing schemes for cloud computing resources in a monopolistic setting, where a single provider sells access to multiple resources and sets prices to optimize revenue [62]. In these models, the provider posts prices for resource usage, users select resource bundles to maximize their individual surplus, and the provider earns revenue from the resulting allocation. Mechanisms such as CloudPack [32] extend this approach by incorporating workload flexibility, using concepts like Shapley values for fair cost distribution among colocated jobs. Similarly, works such as [48] address resource contention through dynamic pricing strategies that adjust costs based on runtime slowdowns.
While structurally similar, the blockchain context introduces an additional constraint. In addition to objectives like maximizing provider revenue, it emphasizes maximizing throughput by ensuring transactions are processed efficiently, minimizing their collective execution time within a block. Moreover, blockchain transactions explicitly declare access to distinct storage keys, introducing contention through overlapping access patterns. This adds complexity compared to cloud models, which typically abstract resource demands as scalar quantities and address contention at a more aggregated level.
Capturing the effects of storage key-level conflicts on parallel execution in blockchains requires fee computation mechanisms that are sensitive to these fine-grained interactions. Thus, while conceptually addressing similar challenges, the blockchain setting imposes unique constraints that motivate the need for specialized GCMs as explored in this work.
8 Conclusion
@@
In this work, we addressed this gap by introducing a framework with two key components: a GCM, which measures the execution-related load a transaction imposes on the network in units of gas, and a TFM, which determines the cost associated with each unit of gas. We then formalized the desired properties for the GCM in such a fee market. After outlining these desiderata, we evaluated various mechanisms against them and identified two strong candidates through this analysis: the weighted area GCM for the class of mechanisms with Easy Gas Estimation, and the time-proportional makespan GCM for the class without it.
@@
Appendix A Proofs Omitted From Section˜3
@@
Proof A.1.
Properties˜1 and 2 are special cases of Property˜3, corresponding to the scenarios where $t_{1}=t_{2}$ and $K_{1}=K_{2}$, respectively. Therefore, the $(\Rightarrow)$ direction holds. To prove the $(\Leftarrow)$ direction, assume Properties˜1 and 2 hold and consider a set of transactions $T$ and two transactions $\textit{tx}_{1}\simeq(t_{1},K_{1})$ and $\textit{tx}_{2}\simeq(t_{2},K_{2})$, both not in $T$, such that $t_{1}\leq t_{2}$ and $K_{1}\subseteq K_{2}$. Let $\textit{tx}_{3}$ be another transaction not in $T$ such that $\textit{tx}_{3}\simeq(t_{1},K_{2})$. By first applying Property˜1 and then Property˜2, we obtain:
@@
This establishes the conclusion.
@@
Proof A.2.
The strict variants of Properties˜1 and 2 are special cases of the strict variant of Property˜3, corresponding to the scenarios where $t_{1}=t_{2}$ and $K_{1}=K_{2}$, respectively. Therefore, the $(\Rightarrow)$ direction holds. To prove the $(\Leftarrow)$ direction, assume Properties˜1 and 2 hold strictly and consider a set of transactions $T$ and two transactions $\textit{tx}_{1}\simeq(t_{1},K_{1})$ and $\textit{tx}_{2}\simeq(t_{2},K_{2})$, both not in $T$, such that $t_{1}\leq t_{2}$ and $K_{1}\subseteq K_{2}$, and additionally $\textit{tx}_{1}\not\simeq\textit{tx}_{2}$ (i.e., at least one of the previous holds strictly). Let $\textit{tx}_{3}$ be another transaction not in $T$ such that $\textit{tx}_{3}\simeq(t_{1},K_{2})$. By first applying Property˜1 and then Property˜2 (their non-strict versions), we obtain:
$\textsf{gas}_{T\cup\{\textit{tx}_{1}\}}(\textit{tx}_{1})\leq\textsf{gas}_{T\cup\{\textit{tx}_{3}\}}(\textit{tx}_{3})\leq\textsf{gas}_{T\cup\{\textit{tx}_{2}\}}(\textit{tx}_{2}).$ (5)
From this, we get that $\textsf{gas}_{T\cup{\textit{tx}_{1}}}(\textit{tx}_{1})\leq\textsf{gas}_{T\cup{\textit{tx}_{2}}}(\textit{tx}_{2})$, so it remains to rule out the equality case. Assume for a contradiction that $\textsf{gas}_{T\cup\{\textit{tx}_{1}\}}(\textit{tx}_{1})=\textsf{gas}_{T\cup\{\textit{tx}_{2}\}}(\textit{tx}_{2})$, from which the two inequalities hold with equality in Equation˜5. Because Properties˜1 and 2 hold strictly, this can only be the case if $\textit{tx}_{1}\simeq\textit{tx}_{3}\simeq\textit{tx}_{2}$, contradicting the assumption that $\textit{tx}_{1}\not\simeq\textit{tx}_{2}$.
Appendix B Proofs Omitted From Section˜4